OpenStudy (anonymous):

Solve for x. y =x/3 + 8

OpenStudy (bradely):

y =x/3 + 8 Subtract 8 both sides y-8 = x/3 x=3(y-8) Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

thanks :)